# Semi-Supervised Clustering of Sparse Graphs: Crossing the Information-Theoretic Threshold

### 08 Jan 2024 v3

@@ 3.2 Locally Tree-Like Structure
Proceeding to the $t$-neighbors, we need to understand a bit better the structure of a small neighborhood in the SBM. The neighborhoods in a sparse network locally have no loops. So they have a nice tree-like structure. Moreover, the labels also obey some random broadcasting processes on trees.
A broadcasting process transmit the information from the root of a tree to all the nodes. At each level, nodes inherit the information from its parent. But error could happen with certain amount of probability. Usually the edges are assumed to be included according to the same rule and work independently. It was firstly considered in genetics (Cavender, 1978) since it perfectly describes the propagation of a gene from ancestor to descendants. It can also be interpreted as a communication network that pass out the information from the root. So such processes were intensively studied in information theory and statistical physics (Spitzer, 1975; Higuchi, 1977; Bleher et al., 1995). In particular, we are interested in the following Markov process since it can be identified with the labeling process of a small neighborhood in SBM.
@@
Let $T$ be an infinite rooted tree with root $v$. Given a number $10\leq\epsilon<$ 1 and the offspring rate $d>0$, we define a random labelling $\tau\in\{1,-1\}^{T}$. First, draw $\tau_{v}$ uniformly in $\{1,-1\}$. Then recursively construct the labelling as follows.
@@
Hence, for fixed $t\in\mathbb{N}$, $t\leq R$ and any $v\not\in\mathcal{R}$, we can denote the label of a vertex in $v$’s $t$-neighborhood as $Y_{i}^{(t)}\coloneqq\Pi_{k=1}^{t}\prescript{k}{}{Y_{u}}$, where $\{\prescript{k}{}{Y_{u}}\}_{k=1}^{t}$ are independent copies of $Y_{u}$. Then we have $\operatorname{E}(Y_{i}^{(t)})=(\frac{a-b}{2n})^{t}$ and $\operatorname{E}((Y_{i}^{(t)})^{2})=(\frac{a+b}{2n})^{t}$. Moreover, $\{Y_{i}^{(t)}\}$’s are essentially independent. Therefore, the census of $v$’s revealed $t$-neighbors can be written as
@@
$\tilde{\Delta}_{t}(v)\to\mathcal{N}(\rho(\frac{a-b}{2})^{t},\rho(\frac{a+b}{2})^{t})\qquad\text{as~{}}n\to\infty$ (24)
@@
$\operatorname{P}(\text{sgn}(\tilde{\Delta}_{1}(v))=x_{v})>\frac{1}{2}+\frac{1}{3}\sqrt{\rho~{}\text{SNR}}$ (27)
@@
$\text{Overlap}(x_{\mathcal{R}^{\complement}},\hat{x}_{\mathcal{R}^{\complement}})=\frac{1}{n-m}|\langle x_{\mathcal{R}^{\complement}},\hat{x}_{\mathcal{R}^{\complement}}\rangle|$ (29)
@@
$>\frac{2}{3}\sqrt{\rho~{}\text{SNR}}$ (32)
The expected overlap is not vanishing which suggests the weak recovery is solvable for any SNR. But it is technically impractical to rigorously describe the limit distribution of our census estimator without blurring this edge out. From Figure 3, we can see that our calculation is close to the expectation. But the convergence rate depends on $\rho$. In particular, when both SNR and $\rho$ are small, the asymptotic behavior of our algorithm remains unclear. Hence we go through a direct analysis to establish the desired result.
Figure 3: The simulation result of $\operatorname{\mathcal{G}}(3000,5/3000,2/3000)$, $\text{SNR}\approx 0.64$. Solid curves stand for the average overlaps of the t-neighbors census method ($t$ = 1, 2, and 3) on 60 independent realizations of the random graph. Shaded area represents the standard error band of 1-neighbors census. The dashed curve stands for the asymptotic lower bound we conclude from our calculation.
@@ 3.3 Majority of 1-Neighbors
$A_{ij}\sim\text{Bernoulli}\left(\frac{a}{n}\right)$ $\quad\left(i\leq\frac{n}{2}~{}\text{and}~{}j\leq\frac{n}{2}\right)$ $~{}\text{or}~{}\left(i\geq\frac{n}{2}~{}\text{and}~{}j\geq\frac{n}{2}\right)$ (33)
$A_{ij}\sim\text{Bernoulli}\left(\frac{b}{n}\right)$ $\quad i\geq\frac{n}{2}~{}$ $\text{and}~{}j\leq\frac{n}{2}$ (34)
@@
$x_{i}=\begin{cases}1,\quad&i=1,2,\dots,\frac{n}{2},\\ -1,\quad&i=\frac{n}{2},\frac{n}{2}+1,\dots,n,\end{cases}\quad\tilde{x}_{i}=\begin{cases}1,\quad&i=1,2,\dots,\frac{m}{2},\\ -1,\quad&i=\frac{n}{2},\frac{n}{2}+1,\dots,\frac{n+m}{2},\\ 0,\quad&\text{otherwise.}\end{cases}$ (35)
@@
$\tilde{\Delta}_{1}(i)=\langle A[i,\mathrel{\mathop{:}}],\tilde{x}\rangle=\sum_{j\mathrel{\mathop{:}}\tilde{x}_{j}\neq 0}A_{ij}\tilde{x}_{j}=\sum_{j\mathrel{\mathop{:}}\tilde{x}_{j}\neq 0}A_{ji}\tilde{x}_{j}$ (36)
Therefore, $\{\tilde{\Delta}_{1}(i)\}$’s are independent for all $i\mathrel{\mathop{:}}\tilde{x}_{i}=0$ since they have no common term. Notice it is not the case for all $i\in[n]$. But we only need to predict the unrevealed labels, hence the independence.
@@
$\hat{x}_{i}=\begin{cases}\tilde{x}_{i}\quad&\text{if}~{}\tilde{x}_{i}\neq 0\\ \text{sgn}^{*}(\tilde{\Delta}_{1}(i))\quad&\text{if}~{}\tilde{x}_{i}=0\end{cases}$ (37)
@@
$\operatorname{P}(\text{sgn}^{*}(\tilde{\Delta}_{1}(i))=1|\tilde{\Delta}_{1}(i)=0)=\operatorname{P}(\text{sgn}^{*}(\tilde{\Delta}_{1}(i))=-1|\tilde{\Delta}_{1}(i)=0)=\frac{1}{2}$ (38)
@@
where $\{B_{i}\}$’s are i.i.d. Bernoulli(1/2) random variables, independent of all other random objects. Notice that it is only introduced for analysis purpose and is equivalent to the conventional sign function in practice.
Suppose $(G,x)$ is an Erdős–Rényi random graph with revealed label $\tilde{x}$, any estimator can only have vanishing correlation with the true label among the unrevealed vertices. So the semi-supervised weak recovery problem on SBM requires finding an estimator such that the correlation restricted on the unrevealed part is non-vanishing. Formally, we want to show that
$\operatorname{P}\left(\text{Overlap}(x|_{\tilde{x}_{i}=0},\hat{x}|_{\tilde{x}_{i}=0})\geq\Omega(1)\right)=1-o(1)$ (40)
As discussed in Section 2.1, we start off with a critical result scrutinizing binomial variable sequences. It gives us an edge over direct analysis via a Berry–Esseen-type inequality, which usually assumes distribution of individual random variable in the sequence independent of $n$.
@@
$\operatorname{P}(X>Y)-\operatorname{P}(X<Y)\geq\delta$ (41)
@@
$=\sum_{x=1}^{n}\sum_{y=0}^{x-1}\left[\binom{n}{x}\left(\frac{a}{n}\right)^{x}\left(1-\frac{a}{n}\right)^{n-x}\binom{n}{y}\left(\frac{b}{n}\right)^{y}\left(1-\frac{b}{n}\right)^{n-y}\right]$ (44)
@@
$\phantom{\Delta}=\sum_{x=1}^{n}\binom{n}{x}\left(\frac{ab}{n}\right)^{x}\left(1-\frac{a+b}{n}+\frac{ab}{n^{2}}\right)^{n-x}\\ \Bigg{\{}\sum_{y=0}^{x-1}\binom{n}{y}\frac{1}{n^{y}}\left[\left(\frac{1}{b}-\frac{1}{n}\right)^{x-y}-\left(\frac{1}{a}-\frac{1}{n}\right)^{x-y}\right]\Bigg{\}}$
@@
Also notice that $\binom{n}{m}=\prod_{i=0}^{m-1}\frac{n-i}{m-i}\geq\left(\frac{n}{m}\right)^{m},~{}\forall 1\leq m\leq n$. We have,
@@
We now convert the KL divergence to the total variation distance, which is easier to work with. Let $P_{1}$ and $P_{2}$ be two probability measures defined on the same sample space $\Omega$ and sigma-algebra $\mathcal{F}$. The total variation distance between them is defined as $d_{TV}(P_{1},P_{2})=\sup_{E\in\mathcal{F}}|P_{1}(E)-P_{2}(E)|$. Moreover, in the discrete case, we have following identity $d_{TV}(P_{1},P_{2})=\frac{1}{2}\|P_{1}-P_{2}\|_{1}=\sum_{\omega\in\Omega}\frac{1}{2}\|P_{1}(\omega)-P_{2}(\omega)\|$. It is related to the KL divergence through Pinsker’s inequality (see, eg. (Tsybakov, 2009), Chapter 3). For completeness, we include an elementary proof of the Bernoulli special case that is sufficient for our usage later.
@@
$D(x\|y)=x\ln\frac{x}{y}+(1-x)\ln(\frac{1-x}{1-y})$ (50)
$2(d_{TV}(P_{1},P_{2}))^{2}=\frac{1}{2}\|P_{1}-P_{2}\|_{1}^{2}=2(x-y)^{2}$ (51)
@@
$\tilde{\Delta}_{1}(i)=\sum_{j\mathrel{\mathop{:}}\tilde{x}_{j}\neq 0}A_{ij}\tilde{x}_{j}=\left(\sum_{\rho\frac{n}{2}<j\leq\frac{n}{2}}A_{ij}\right)-\left(\sum_{(1+\rho)\frac{n}{2}<j\leq n}A_{ij}\right)$ (54)
@@
for sufficiently large $n$. Also notice that
$\operatorname{P}(\text{sgn}(\tilde{\Delta}_{1}(i))=-x_{i})=1-\operatorname{P}(\text{sgn}(\tilde{\Delta}_{1}(i))=x_{i})-\operatorname{P}(\tilde{\Delta}_{1}(i)=0)$ (56)
@@
$\operatorname{P}(\text{sgn}(\tilde{\Delta}_{1}(i))=x_{i})\geq\frac{1+\delta}{2}-\frac{1}{2}\operatorname{P}(\tilde{\Delta}_{1}(i)=0)$ (57)
@@
$\geq\frac{1}{2}+\frac{\delta}{2}$ (60)
Since $\{\hat{x}_{i}\}$’s are independent for all unrevealed vertices as $\{\tilde{\Delta}_{1}(i)\}$’s and $\operatorname{E}\left[\frac{\hat{x}_{i}x_{i}+1}{2}\right]=\operatorname{P}(\hat{x}_{i}=x_{i})$, Lemma 17 and Lemma 18 give us that
$\operatorname{P}\left(\frac{1}{(1-\rho)n}\sum_{i\mathrel{\mathop{:}}\tilde{x}_{i}=0}\frac{\hat{x}_{i}x_{i}+1}{2}\leq\frac{1}{2}+\frac{\delta}{2}-\epsilon\right)\leq\mathrm{e}^{-2\epsilon^{2}(1-\rho)n}$ (61)
@@
$\operatorname{P}\left(\text{Overlap}(x|_{\tilde{x}_{i}=0},\hat{x}|_{\tilde{x}_{i}=0})\geq\frac{\delta}{2}\right)\geq 1-\mathrm{e}^{-\frac{\delta^{2}(1-\rho)n}{8}}$ (62)
@@
$P_{n}^{(0)}(E_{n})\rightarrow 0\quad$ (Law of large number) (63)
$P_{n}^{(1)}(E_{n})\nrightarrow 0\quad$ (Bounded from below) (64)

### 28 Feb 2024 v4

@@ 3.2 Locally Tree-Like Structure
Proceeding to the $t$-neighbors, we need to first understand the structure of a small neighborhood in the SBM. The neighborhoods in a sparse network locally have no loops. So they have a nice tree-like structure. Moreover, the labels also obey some random broadcasting processes on trees.
A broadcasting process transmits the information from the root of a tree to all the nodes. At each level, nodes inherit the information from its parent. Meanwhile, errors could happen with a certain amount of probability. Usually, the edges are assumed to be included according to the same rule and work independently. It was first considered in genetics (Cavender, 1978) since it perfectly describes the propagation of a gene from ancestor to descendants. It can also be interpreted as a communication network that passes out the information from the root. Such processes were intensively studied in information theory and statistical physics (Spitzer, 1975; Higuchi, 1977; Bleher et al., 1995). In particular, we are interested in the following Markov process since it can be identified with the labeling process of a small neighborhood in SBM.
@@
Let $T$ be an infinite rooted tree with root $v$. Given a number $10\leq\epsilon<$ 1 and the offspring rate $d>0$, we define a random labeling $\tau\in\{1,-1\}^{T}$. First, draw $\tau_{v}$ uniformly in $\{1,-1\}$. Then, recursively construct the labeling as follows:
@@
Hence, for fixed $t\in\mathbb{N}$, $t\leq R$, and any $v\not\in\mathcal{R}$, we can denote the label of a vertex in $v$’s $t$-neighborhood as $Y_{i}^{(t)}\coloneqq\Pi_{k=1}^{t}\prescript{k}{}{Y_{u}}$, where $\{\prescript{k}{}{Y_{u}}\}_{k=1}^{t}$ are independent copies of $Y_{u}$. Then we have $\operatorname{E}(Y_{i}^{(t)})=(\frac{a-b}{2n})^{t}$ and $\operatorname{E}((Y_{i}^{(t)})^{2})=(\frac{a+b}{2n})^{t}$. Moreover, $\{Y_{i}^{(t)}\}$’s are essentially independent. Therefore, the census of $v$’s revealed $t$-neighbors can be written as
@@
$\tilde{\Delta}_{t}(v)\to\mathcal{N}(\rho(\frac{a-b}{2})^{t},\rho(\frac{a+b}{2})^{t}),\quad\text{as~{}}n\to\infty.$ (24)
@@
$\operatorname{P}(\text{sgn}(\tilde{\Delta}_{1}(v))=x_{v})>\frac{1}{2}+\frac{1}{3}\sqrt{\rho~{}\text{SNR}}.$ (27)
@@
$\text{Overlap}(x_{\mathcal{R}^{\complement}},\hat{x}_{\mathcal{R}^{\complement}})=\frac{1}{n-m}|\langle x_{\mathcal{R}^{\complement}},\hat{x}_{\mathcal{R}^{\complement}}\rangle|.$ (29)
@@
$>\frac{2}{3}\sqrt{\rho~{}\text{SNR}}.$ (32)
The expected overlap is not vanishing which suggests that weak recovery is solvable for any SNR. But it is technically impractical to rigorously describe the limit distribution of our census estimator without blurring this edge out. From Figure 3, we can see that our calculation is close to the expectation. But the convergence rate depends on $\rho$. In particular, when both SNR and $\rho$ are small, the asymptotic behavior of our algorithm remains unclear. Hence, we go through a direct analysis to establish the desired result.
Figure 3: The simulation result of $\operatorname{\mathcal{G}}(3000,5/3000,2/3000)$, $\text{SNR}\approx 0.64$. Solid curves stand for the average overlaps of the t-neighbors census method ($t$ = 1, 2, and 3) on 60 independent realizations of the random graph. The shaded area represents the standard error band of the 1-neighbors census. The dashed curve stands for the asymptotic lower bound we conclude from our calculation.
@@ 3.3 Majority of 1-Neighbors
$A_{ij}\sim\text{Bernoulli}\left(\frac{a}{n}\right)$ $\quad\left(i\leq\frac{n}{2}~{}\text{and}~{}j\leq\frac{n}{2}\right)$ $~{}\text{or}~{}\left(i\geq\frac{n}{2}~{}\text{and}~{}j\geq\frac{n}{2}\right),$ (33)
$A_{ij}\sim\text{Bernoulli}\left(\frac{b}{n}\right)$ $\quad i\geq\frac{n}{2}~{}$ $\text{and}~{}j\leq\frac{n}{2}.$ (34)
@@
$x_{i}=\begin{cases}1\quad&i=1,2,\dots,\frac{n}{2},\\ -1\quad&i=\frac{n}{2},\frac{n}{2}+1,\dots,n;\end{cases}\quad\tilde{x}_{i}=\begin{cases}1\quad&i=1,2,\dots,\frac{m}{2},\\ -1\quad&i=\frac{n}{2},\frac{n}{2}+1,\dots,\frac{n+m}{2},\\ 0\quad&\text{otherwise.}\end{cases}$ (35)
@@
$\tilde{\Delta}_{1}(i)=\langle A[i,\mathrel{\mathop{:}}],\tilde{x}\rangle=\sum_{j\mathrel{\mathop{:}}\tilde{x}_{j}\neq 0}A_{ij}\tilde{x}_{j}=\sum_{j\mathrel{\mathop{:}}\tilde{x}_{j}\neq 0}A_{ji}\tilde{x}_{j}.$ (36)
Therefore, $\{\tilde{\Delta}_{1}(i)\}$’s are independent for all $i\mathrel{\mathop{:}}\tilde{x}_{i}=0$ since they have no common term. Note that this is not the case for all $i\in[n]$. However, we only need to predict the unrevealed labels, hence the independence holds.
@@
$\hat{x}_{i}=\begin{cases}\tilde{x}_{i}\quad&\text{if}~{}\tilde{x}_{i}\neq 0,\\ \text{sgn}^{*}(\tilde{\Delta}_{1}(i))\quad&\text{if}~{}\tilde{x}_{i}=0.\end{cases}$ (37)
@@
$\operatorname{P}(\text{sgn}^{*}(\tilde{\Delta}_{1}(i))=1|\tilde{\Delta}_{1}(i)=0)=\operatorname{P}(\text{sgn}^{*}(\tilde{\Delta}_{1}(i))=-1|\tilde{\Delta}_{1}(i)=0)=\frac{1}{2}.$ (38)
@@
where $\{B_{i}\}$’s are i.i.d. Bernoulli(1/2) random variables, independent of all other random objects. Note that it is only introduced for analysis purposes and is equivalent to the conventional sign function in practice.
Suppose $(G,x)$ is an Erdős–Rényi random graph with revealed label $\tilde{x}$, any estimator can only have a vanishing correlation with the true label among the unrevealed vertices. So the semi-supervised weak recovery problem on SBM requires finding an estimator such that the correlation restricted on the unrevealed part is non-vanishing. Formally, we want to show that
$\operatorname{P}\left(\text{Overlap}(x|_{\tilde{x}_{i}=0},\hat{x}|_{\tilde{x}_{i}=0})\geq\Omega(1)\right)=1-o(1).$ (40)
As discussed in Section 2.1, we start with a critical result scrutinizing binomial variable sequences. It gives us an edge over direct analysis via a Berry–Esseen-type inequality, which usually assumes the distribution of individual random variables in the sequence independent of $n$.
@@
$\operatorname{P}(X>Y)-\operatorname{P}(X<Y)\geq\delta.$ (41)
@@
$=\sum_{x=1}^{n}\sum_{y=0}^{x-1}\left[\binom{n}{x}\left(\frac{a}{n}\right)^{x}\left(1-\frac{a}{n}\right)^{n-x}\binom{n}{y}\left(\frac{b}{n}\right)^{y}\left(1-\frac{b}{n}\right)^{n-y}\right].$ (44)
@@
$\phantom{\Delta}=\sum_{x=1}^{n}\binom{n}{x}\left(\frac{ab}{n}\right)^{x}\left(1-\frac{a+b}{n}+\frac{ab}{n^{2}}\right)^{n-x}\\ \Bigg{\{}\sum_{y=0}^{x-1}\binom{n}{y}\frac{1}{n^{y}}\left[\left(\frac{1}{b}-\frac{1}{n}\right)^{x-y}-\left(\frac{1}{a}-\frac{1}{n}\right)^{x-y}\right]\Bigg{\}}.$
@@
Also notice that $\binom{n}{m}=\prod_{i=0}^{m-1}\frac{n-i}{m-i}\geq\left(\frac{n}{m}\right)^{m},~{}\forall 1\leq m\leq n$. We have
@@
We now convert the KL divergence to the total variation distance, which is easier to work with. Let $P_{1}$ and $P_{2}$ be two probability measures defined on the same sample space $\Omega$ and sigma-algebra $\mathcal{F}$. The total variation distance between them is defined as $d_{TV}(P_{1},P_{2})=\sup_{E\in\mathcal{F}}|P_{1}(E)-P_{2}(E)|$. Moreover, in the discrete case, we have the following identity $d_{TV}(P_{1},P_{2})=\frac{1}{2}\|P_{1}-P_{2}\|_{1}=\sum_{\omega\in\Omega}\frac{1}{2}\|P_{1}(\omega)-P_{2}(\omega)\|$. It is related to the KL divergence through Pinsker’s inequality (see, e.g., (Tsybakov, 2009), Chapter 3). For completeness, we include an elementary proof of the Bernoulli special case that is sufficient for our usage later.
@@
$D(x\|y)=x\ln\frac{x}{y}+(1-x)\ln(\frac{1-x}{1-y}),$ (50)
$2(d_{TV}(P_{1},P_{2}))^{2}=\frac{1}{2}\|P_{1}-P_{2}\|_{1}^{2}=2(x-y)^{2}.$ (51)
@@
$\tilde{\Delta}_{1}(i)=\sum_{j\mathrel{\mathop{:}}\tilde{x}_{j}\neq 0}A_{ij}\tilde{x}_{j}=\left(\sum_{\rho\frac{n}{2}<j\leq\frac{n}{2}}A_{ij}\right)-\left(\sum_{(1+\rho)\frac{n}{2}<j\leq n}A_{ij}\right).$ (54)
@@
for sufficiently large $n$. Also, notice that
$\operatorname{P}(\text{sgn}(\tilde{\Delta}_{1}(i))=-x_{i})=1-\operatorname{P}(\text{sgn}(\tilde{\Delta}_{1}(i))=x_{i})-\operatorname{P}(\tilde{\Delta}_{1}(i)=0).$ (56)
@@
$\operatorname{P}(\text{sgn}(\tilde{\Delta}_{1}(i))=x_{i})\geq\frac{1+\delta}{2}-\frac{1}{2}\operatorname{P}(\tilde{\Delta}_{1}(i)=0).$ (57)
@@
$\geq\frac{1}{2}+\frac{\delta}{2}.$ (60)
Since the $\{\hat{x}_{i}\}$’s are independent for all unrevealed vertices as $\{\tilde{\Delta}_{1}(i)\}$’s and $\operatorname{E}\left[\frac{\hat{x}_{i}x_{i}+1}{2}\right]=\operatorname{P}(\hat{x}_{i}=x_{i})$, Lemma 17 and Lemma 18 give us that
$\operatorname{P}\left(\frac{1}{(1-\rho)n}\sum_{i\mathrel{\mathop{:}}\tilde{x}_{i}=0}\frac{\hat{x}_{i}x_{i}+1}{2}\leq\frac{1}{2}+\frac{\delta}{2}-\epsilon\right)\leq\mathrm{e}^{-2\epsilon^{2}(1-\rho)n}.$ (61)
@@
$\operatorname{P}\left(\text{Overlap}(x|_{\tilde{x}_{i}=0},\hat{x}|_{\tilde{x}_{i}=0})\geq\frac{\delta}{2}\right)\geq 1-\mathrm{e}^{-\frac{\delta^{2}(1-\rho)n}{8}}.$ (62)
@@
[…]
Recall that distinguishability and weak recovery are equivalent in the sense that they share the same phase transition threshold. In the semi-supervised setting, it is straightforward that the weak recovery implies distinguishability. So, Theorem 11 suggests that in this case, SBM and ERM are always distinguishable, which is also equivalent to Corollary 19. Hence, the two problems share the same solvable region as well.